9. Evaluate the function (f -9) (-2) for: f (x) = x2 +4, and g(x) = x-5 . 0-5 O 8 O 11 0-15 0 15
Answer #1

Given fins=art 4 =) +-2)=(25+4 te-2 )=4+4=8 [put n=- 2 7 Giren glas=x-5 = g(-2)=-2-5 r. (-9) (-2) = +-2)-g(-2) [by linearity